1. Plot these data as electromagnetic waves on a graph and find the following:


a. Wavelength


b. Amplitude



1.1 Highest position (y)= 12m


Lowest position (y)= -4m


Cycles= 6


Distance (x) = 32m



1.2 Highest position (y)= 15m


Lowest position (y)= -3m


Cycles= 4


Distance (x) = 25m




1
Expert's answer
2022-05-03T01:29:09-0400

1.1Wavelength =Distance/Cycles=32/6=5.33 m

Amplitude =Highest position - Lowest position=12-(-4)=16 m

1.2 Wavelength =Distance/Cycles=25/4=6.25 m


Amplitude =Highest position - Lowest position=15-(-3)=18 m